# Neat: Nonlinear Parameter-efficient Adaptation of Pre-trained Models

### v2

@@ Abstract
Fine-tuning pre-trained models often yields state-of-the-art performance but is computationally expensive when updating all parameters. Parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA), address this by freezing pre-trained weights and introducing low-rank matrices. However, because LoRA relies on low-rank decomposition, it struggles to capture complex nonlinear dynamics and optimal optimization trajectories, resulting in a performance gap relative to full fine-tuning and inefficient parameter utilization. To overcome these issues, we propose Neat, a nonlinear PEFT approach that employs a lightweight neural network to learn a nonlinear transformation of the pre-trained weights, thereby better approximating cumulative weight updates. Our theoretical analysis shows that Neat achieves greater efficiency than LoRA while maintaining equivalent expressivity. Extensive experiments on four benchmarks and over twenty datasets demonstrate that Neat significantly outperforms state-of-the-art baselines in both vision and text tasks.
$\bullet$\$$\bullet$\$footnotetext: These authors contributed equally to this work
@@ 1 Introduction
Pre-trained models, trained on extensive and diverse general-domain corpora, demonstrate exceptional generalization capabilities, benefiting a range of fundamental tasks, such as natural language understanding (Devlin, 2018; Liu, 2019), natural language generation (Touvron et al., 2023; AI@Meta, 2024), and image classification (Dosovitskiy et al., 2020). In order to adapt pre-trained models to specific downstream tasks, fine-tuning is typically employed. However, due to the large number of parameters in pre-trained models, full fine-tuning requires significant computational resources and incurs substantial memory overhead (Qin et al., 2024).
To address this challenge, various parameter-efficient fine-tuning (PEFT) techniques (Ding et al., 2023; Han et al., 2024) have been developed, enabling pre-trained models to be fine-tuned in resource-constrained environments (Lin et al., 2024). PEFT methods reduce the memory overhead of fine-tuning by introducing a small set of learnable parameters, updating only these lightweight components. These approaches allow pre-trained models to effectively adapt to new tasks while minimizing resource consumption. Among PEFT techniques, the Low-Rank Adaptation (LoRA) family (Hu et al., 2021a; Liu et al., 2024; Song et al., 2024; Büyükakyüz, 2024; Zhao et al., 2024) is widely regarded as one of the most effective and popular approaches due to its minimal architectural modifications, high efficiency, and strong performance. Instead of updating pre-trained model weight directly, LoRA instead introduces two learnable low-rank matrices for it, and approximate weight updates through their product. Since these low-rank matrices are much smaller than the original pre-trained weights, LoRA significantly reduces the memory overhead during fine-tuning.
Despite its widespread success, LoRA has limitations, particularly in capturing complex relationships in weight updates. By decomposing weight updates into low-rank matrices, LoRA effectively reduces the fine-tuning parameter space, but this comes at the cost of failing to capture complex patterns that are critical for many downstream tasks (Pan et al., 2024). Specifically, LoRA approximation often struggles to model the intricate optimization trajectories required for high performance, especially when the rank of the low-rank matrices is small. Consequently, LoRA-based methods often require higher ranks to close the gap between their performance and that of full fine-tuning, which increases the number of parameters and undermines the efficiency gains they were designed to achieve.
@@
To overcome these limitations, we propose a novel nonlinear parameter-efficient adaptation method, Neat, which incorporates a lightweight neural network into the adaptation process. Unlike LoRA, which approximates weight updates linearly through low-rank decomposition, Neat models cumulative weight updates as explicit functions of the pre-trained model’s original weights. This enables Neat to capture complex, non-linear patterns in the weight space, improving adaptation performance without increasing the number of parameters. The key innovation in Neat lies in introducing a neural network that transforms the pre-trained weights, approximating the updates with minimal additional computation. This nonlinear transformation enhances the expressiveness of the parameter updates while maintaining the efficiency. Importantly, this architecture facilitates a more efficient exploration of the optimization landscape, leading to better task adaptation, particularly in cases where linear methods like LoRA would require much larger ranks to achieve competitive results. We theoretically demonstrate that Neat can achieve the same or greater expressivity than LoRA with fewer parameters.
The contributions are summarized as follows: 1) We propose Neat, a novel parameter-efficient fine-tuning method that leverages nonlinear transformations, effectively capturing more complex weight updates. To the best of our knowledge, this is the first work to introduce nonlinear adaptation for LoRA-based PEFT methods; 2) The proposed Neat enhances model performance while maintaining the efficiency. We theoretically show that Neat can achieve a possibly improved parameter efficiency compared to LoRA; 3) We conduct extensive experiments on four benchmarks covering over twenty datasets. The experiments show that the proposed Neat can outperform baselines on both vision and text tasks.
2 Related Works
@@
In this section, we provide a concise overview of related work on Parameter-Efficient Fine-Tuning (PEFT) methods. PEFT methods aim to reduce the memory overhead of fine-tuning pre-trained models, enabling fine-tuning in resource-constrained environments. According to Han et al. (2024), PEFT methods can be categorized into: 1) Additive PEFT methods (Chronopoulou et al., 2023; Edalati et al., 2022; Lester et al., 2021; Wang et al., 2024a; Liu et al., 2022), 2) Selective PEFT methods (Guo et al., 2020; Das et al., 2023; Sung et al., 2021; Ansell et al., 2021; Zaken et al., 2021; Vucetic et al., 2022), 3) Reparameterized PEFT methods (Hu et al., 2021b; Valipour et al., 2022; Zhang et al., 2023; Karimi Mahabadi et al., 2021; Liu et al., 2024; Kopiczko et al., 2023), and 4) Hybrid PEFT methods (Mao et al., 2021; Chen et al., 2023; He et al., 2021; Zhang et al., 2022; Zhou et al., 2024). Among these, Low-Rank Adaptation (LoRA)-based methods, which are representative of reparameterized PEFT approaches, have gained significant attention due to their minimal architectural changes, no additional inference costs, and high efficiency. LoRA (Hu et al., 2021b) introduces two trainable low-rank matrices for each pre-trained model weight to approximate the desired updates of the original model. Extensions of LoRA include DyLoRA (Valipour et al., 2022), which dynamically adjusts the rank of the low-rank matrices during training to optimize for specific tasks; AdaLoRA (Zhang et al., 2023), which adaptively allocates the parameter budget among weight matrices based on their importance scores; and DoRA (Liu et al., 2024), which decomposes the pre-trained weight into magnitude and direction, applying LoRA only for direction updates. Other variants include VeRA (Kopiczko et al., 2023), which introduces shared frozen random matrices across layers to improve efficiency further, and RoseLoRA (Wang et al., 2024b), which employs a row- and column-wise sparse low-rank adaptation mechanism to selectively update the most significant parameters. FourierFT (Gao et al., ) replaces the matrix multiplication in LoRA with a Fourier transform, while PiSSA (Meng et al., 2024) and MiLoRA (Wang et al., 2024c) update the principal and minor singular components of the weight matrix, respectively. However, existing PEFT methods rely on linear transformations to approximate pre-trained weight updates, which struggle to capture the complex relationships inherent in weight updates, leading to a significant performance gap compared to full fine-tuning.
@@ 3 Methodology
In this section, we start with a brief introduction of LoRA. Motivated by a key limitation in LoRA parameter efficiency that roots from LoRA parameterization form, we propose Neat, a novel PEFT method to solve the issue. Notably, Neat is able to achieves better parameter efficiency provably.
@@ 3.1 Preliminary
LoRA (Hu et al., 2021b) assumes that the updates to model weights during the fine-tuning exhibit low-rank properties. Built upon this, LoRA models the incremental update of some weight matrix $\mathbf{W}^{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ in a pre-trained model approximately by the product of two learnable low-rank matrices
@@ 3.2 Inherent Limitation of LoRA Formulation
Specifically, when fully fine-tuning a pre-trained model, the update process of weight $\bm{W}$ is typically performed through an iterative gradient descent:
@@
Nonetheless, LoRA matrices $\mathbf{A}$ and $\mathbf{B}$ are parameterized in a free way without any dependency on $\mathbf{W}^{0}$. While gradient $\nabla_{\bm{A}}\mathcal{L}$ and $\nabla_{\bm{B}}\mathcal{L}$ are implicit functions of $\mathbf{W}^{0}$, making final learned $\mathbf{A}_{t},\mathbf{B}_{t}$ indirectly depends on $\mathbf{W}^{0}$ as well, as will be proved shortly, the lack of explicit dependency still makes LoRA inherently suboptimal for fine-tuning pre-trained models.
3.3 Nonlinear Parameter-efficient Adaptation
Motivated by the above analysis on LoRA’s limitation, we propose to approximate $\Delta\mathbf{W}$ using a lightweight neural network that explicitly takes pre-trained model weight $\mathbf{W}^{0}$ as input and outputs the weight update directly. By doing so, our approach captures more complex and richer transformation of the weights in a more efficient manner. We refer to our method as nonlinear parameter-efficient adaptation method (Neat).
Inspired by the effectiveness of LoRA’s incremental updates paradigm, the proposed Neat also provides incremental update of pre-trained models. However, Neat modifies the forward pass of the model by introducing a dynamic nonlinear weight transformation. Specifically, the modified model’s forward propagation is formulated as:
@@ 3.3 Nonlinear Parameter-efficient Adaptation
To ensure the parameter efficiency of our Neat, the learnable neural network $f(\mathbf{W}^{0};\bm{\theta})$ should be lightweight, i.e., the number of parameters in $f(\mathbf{W}^{0};\bm{\theta})$ should be much fewer than that of the original pre-trained weight $\mathbf{W}^{0}$. Therefore, we parametrize $f(\mathbf{W}^{0};\bm{\theta})$ as a neural network with bottleneck layers. For example, a simple case is $f(\mathbf{W}^{0};\bm{\theta})=\sigma(\mathbf{W}^{0}\bm{\Theta}_{1})\bm{\Theta}_{2}$, where $\bm{\theta}=(\bm{\Theta}_{1},\bm{\Theta}_{2})\in\mathbb{R}^{d_{2}\times r}\times\mathbb{R}^{r\times d_{2}}$ with $r\ll\min(d_{1},d_{2})$, and $\sigma(\cdot)$ is some activation function such as ReLU (Glorot et al., 2011). We can also increase the layers or add activation function for the output of $f(\mathbf{W}^{0};\bm{\theta})$ to enhance the model expressiveness.
@@
where the original pre-trained weight $\mathbf{W}^{0}$ is frozen, and only neural network parameters $\theta$ are updated. The overview of Neat is shown in Fig. 1.
Remark 1.
The benefit of our new formulation lies in two folds. First, our incremental update $f(\mathbf{W}^{0};\bm{\theta})$ is an explicit function of $\mathbf{W}^{0}$, allowing it to capture updates in a more effective way. Second, the neural network-based $f(\mathbf{W}^{0};\bm{\theta})$ allows for dynamic, non-linear weight updates that can capture more complex interactions. These two advantages make Neat a more effective and efficient PEFT method than existing LoRA-based approaches.
Figure 1: Framework of proposed Neat.
@@ 3.4 Theoretical Analysis
We end up this section with a theoretical analysis of the sub-optimality of LoRA parameter efficiency by proving that Neat is capable of achieving equivalent or even higher efficiency certain conditions. Specifically, suppose Neat adopts the following lightweight architecture as mentioned in Sec 3.3
@@
Then below proposition shows that, Neat can achieve the same expressivity offered by LoRA using fewer parameters under certain conditions. Here the expressivity is measured in terms of minimum attainable loss.
@@
In words, Prop 3.1 demonstrates the (approximate) equivalence of LoRA and Neat in terms of their expressivity. Specifically, the minimum attainable loss using rank-$r$ LoRA can be achieved by Neat with $2r$ hidden units, and conversely, the minimum attainable loss using Neat with $r$ hidden units can be achieved rank-$r$ LoRA, provided the invariance assumption holds. This equivalence further implies that the function classes realized by Neat with $O(r)$ hidden dimensions and rank-$r$ LoRA are equivalent in expressivity, as the result holds for any loss functions.
Importantly, this highlights a potential improvement in parameter efficiency by Neat. Namely, Neat with $O(rd_{2})$ parameters maintains the expressivity of LoRA with $r(d_{1}+d_{2})$ parameters. That it to say, Neat offers a significant improvement in parameter efficiency when $d_{2}\ll d_{1}$ ( a condition that widely holds for the down projection matrix of transformers fully-connected layers (Vaswani, 2017; Dosovitskiy et al., 2021) ). In such cases, Neat provably achieves better parameter efficiency than LoRA. The added parameter efficiency can also improve sample efficiency by allowing the model to learn representations with the same or fewer data points.
The invariance assumption in Proposition 3.1 pertains to the pre-trained model, and asserts that the later layers of the model depends solely on the task-relevant feature space. Given that we fine-tune a pre-trained model, the later layers are expected to capture this task-relevant feature space, which is described by the left singular space of $\mathbf{W}^{0}$. In practice, since the later layers primarily rely on this pre-trained feature space, the principal directions of the pre-trained weight matrix, represented by its singular vectors, encode most of the useful features for downstream tasks. This makes the loss largely invariant to changes outside this subspace. The proof is available in Section A.1.
If we consider a sinusoid activation function $\sigma_{\textnormal{p}}(x)=\sin(2\pi x)$, then stronger result that Neat has expressivity (almost) greater than or equal to a LoRA with possibly more parameters can be established without the invariance assumption. We defer the result to the Appendix A.2.

### v3

# PEANuT: Parameter-Efficient Adaptation with Weight-aware Neural Tweakers

@@ Abstract
Fine-tuning large pre-trained foundation models often yields excellent downstream performance but is prohibitively expensive when updating all parameters. Parameter-efficient fine-tuning (PEFT) methods such as LoRA alleviate this by introducing lightweight update modules, yet they commonly rely on weight-agnostic linear approximations, limiting their expressiveness. In this work, we propose PEANuT, a novel PEFT framework that introduces weight-aware neural tweakers, compact neural modules that generate task-adaptive updates conditioned on frozen pre-trained weights. PEANuT provides a flexible yet efficient way to capture complex update patterns without full model tuning. We theoretically show that PEANuT achieves equivalent or greater expressivity than existing linear PEFT methods with comparable or fewer parameters. Extensive experiments across four benchmarks with over twenty datasets demonstrate that PEANuT consistently outperforms strong baselines in both NLP and vision tasks, while maintaining low computational overhead.
Keywords: parameter-efficient fine-tuning, foundation model
= Date: November 24, 2025
[Uncaptioned image] Code Repository: https://github.com/yibozhong/peanut
= Contact: yibozhong657@gmail.com; hjiang2@albany.edu; hwang28@albany.edu
[…]
@@ 1 Introduction
Pre-trained models, trained on large and diverse general-domain corpora, have demonstrated strong generalization capabilities across a variety of tasks, including natural language understanding (devlin2018bert, liu2019roberta, howard2018universal, wu2019enriching, sun2023text, wang2021knowledge, wang2022fedkc), generation (llama2-7b, llama3, xu2025collab, liu2025roserag, wang2024blendfilter, yao2022react, lewis2020retrieval), and vision tasks such as image classification (dosovitskiy2020image, bhojanapalli2021understanding, chen2021crossvit). A common strategy for adapting these models to specific downstream tasks is full fine-tuning. However, due to the massive number of parameters involved, full fine-tuning often leads to significant computational and memory costs (qin2024empirical).
To mitigate these challenges, various parameter-efficient fine-tuning (PEFT) methods (ding2023parameter, han2024parameter) have been developed, enabling pre-trained models to be fine-tuned in resource-constrained environments (lin2024awq). These methods retain most of the pre-trained weights in a frozen state and introduce a small set of trainable components, thereby significantly reducing memory and compute overhead (lin2024awq). Among them, Low-Rank Adaptation (LoRA) (lora, liu2024dora, song2024low, buyukakyuz2024olora, zhao2024galore) is a popular and widely adopted approach due to its simplicity, strong empirical performance, and compatibility with modern architectures.
Instead of updating pre-trained model weight directly, LoRA introduces two learnable low-rank matrices for it, and approximate weight updates through their product. Since the numbers of parameters of these low-rank matrices are much smaller than that of the original pre-trained weights, LoRA significantly reduces the memory overhead during fine-tuning.
Despite its widespread success, LoRA has inherent limitations, particularly in its ability to model complex weight adaptation behaviors. LoRA approximates the weight change with the product of two low-rank matrices. While recent studies have observed that the cumulative weight updates during fine-tuning often exhibit approximately low-rank structure (zhao2024galore), LoRA itself learns these updates from scratch using randomly initialized parameters, without leveraging any prior knowledge from the pre-trained weights. As a result, the optimization process becomes more challenging, especially under low-rank settings where the parameter space is highly constrained and prone to suboptimal local minima (pan2024lisa). Furthermore, due to its linear structure, LoRA may struggle to capture intricate adaptation patterns required by many downstream tasks. To compensate for this limited capacity, LoRA-based methods often resort to increasing the rank of the update matrices, which in turn reduces their parameter efficiency and undermines their original motivation.
@@
To overcome these limitations, we propose a parameter-efficient adaptation method with weight-aware neural tweakers, PEANuT, which incorporates a lightweight neural network, which takes the pre-traiend weight as the input, into the adaptation process. Unlike LoRA, which approximates weight updates linearly through low-rank decomposition, PEANuT models cumulative weight updates as explicit functions of the pre-trained model’s original weights. This enables PEANuT to capture complex, non-linear patterns in the weight space, improving adaptation performance without increasing the number of parameters. The key innovation in PEANuT lies in introducing compact neural networks, neural tweakers, that transforms the pre-trained weights, approximating the updates with minimal additional computation. This nonlinear transformation enhances the expressiveness of the parameter updates while maintaining the efficiency. Importantly, this architecture facilitates a more efficient exploration of the optimization landscape, leading to better task adaptation, particularly in cases where linear methods like LoRA would require much larger ranks to achieve competitive results. We theoretically demonstrate that PEANuT can achieve the same or greater expressivity than LoRA with fewer parameters.
The contributions are summarized as follows:
  • We propose PEANuT, a new PEFT method that introduces weight-aware neural tweakers to generate adaptive update signals. The method enables efficient and flexible adaptation beyond linear constraints. To the best of our knowledge, this is the first work to introduce nonlinear adaptation for LoRA-based PEFT methods.
  • The proposed PEANuT enhances model performance while maintaining the efficiency. We theoretically show that PEANuT can achieve a possibly improved parameter efficiency compared to LoRA.
  • We conduct extensive experiments on four benchmarks covering over twenty datasets. The experiments show that the proposed PEANuT can outperform baselines on both vision and text tasks.
2 Related Work
In this section, we provide a concise overview of related work on Parameter-Efficient Fine-Tuning (PEFT) methods. PEFT methods aim to reduce the memory overhead of fine-tuning pre-trained models, enabling fine-tuning in resource-constrained environments. According to han2024parameter, PEFT methods can be categorized into: 1) Additive PEFT methods (chronopoulou2023adaptersoup, edalati2022krona, lester2021power, wang2024universality, liu2022few), 2) Selective PEFT methods (guo2020parameter, das2023unified, sung2021training, ansell2021composable, zaken2021bitfit, vucetic2022efficient, chen2024large, miao2025coeff, chen2025sparse), 3) Reparameterized PEFT methods (hu2021lora, valipour2022dylora, zhang2023adalora, karimi2021compacter, liu2024dora, kopiczko2023vera), and 4) Hybrid PEFT methods (mao2021unipelt, chen2023parameter, he2021towards, zhang2022neural, zhou2024autopeft). Additive PEFT methods (chronopoulou2023adaptersoup, edalati2022krona, lester2021power, wang2024universality, liu2022few) introduces a small set of additional trainable parameters strategically placed within the model. One of the most prominent additive PEFT approaches is Adapter (chronopoulou2023adaptersoup, edalati2022krona, zhao2022tiny), which involves inserting small adapter layers between pre-trained weight blocks. Prompt Tuning (wang2024universality, lester2021power, vu2021spot, li2021prefix) is another technique, where learnable vectors, or "soft prompts," are prepended to the input sequence without modifying the model’s weights. This method is particularly effective for large-scale models and has inspired variants such as Prefix Tuning (li2021prefix). Selective PEFT focuses on optimizing the fine-tuning process by selectively adjusting a subset of the model’s parameters rather than introducing additional ones. For instance, Diff Pruning (guo2020parameter) uses a learnable binary mask to select parameters for fine-tuning. Similarly, FishMask (sung2021training) and Fish-Dip (das2023unified) leverage Fisher information to determine parameter importance and identify the most crucial ones for updates. Additionally, BitFit (zaken2021bitfit) fine-tunes only the bias terms in the model, significantly reducing the number of trainable parameters. Hybrid PEFT methods aim to combine the strengths of various existing PEFT techniques to enhance model performance across diverse tasks. UniPELT (mao2021unipelt) integrates LoRA, prefix-tuning, and adapters within each Transformer block, employing a gating mechanism to determine which module should be active during fine-tuning. S4 (chen2023parameter) further explores the design space by partitioning layers into groups and assigning different PEFT methods to each group. Additionally, NOAH (zhang2022neural) and AUTOPEFT (zhou2024autopeft) leverage neural architecture search (NAS) to automatically discover optimal combinations of PEFT techniques tailored to specific tasks.
@@ 2 Related Work
Reparameterized PEFT methods are most close to our proposed method. Low-Rank Adaptation (LoRA)-based methods, which are representative of reparameterized PEFT approaches, have gained significant attention due to their minimal architectural changes, no additional inference costs, and high efficiency. LoRA (hu2021lora) introduces two trainable low-rank matrices for each pre-trained model weight to approximate the desired updates of the original model. Extensions of LoRA include DyLoRA (valipour2022dylora), which dynamically adjusts the rank of the low-rank matrices during training to optimize for specific tasks; AdaLoRA (zhang2023adalora), which adaptively allocates the parameter budget among weight matrices based on their importance scores; and DoRA (liu2024dora), which decomposes the pre-trained weight into magnitude and direction, applying LoRA only for direction updates. Other variants include VeRA (kopiczko2023vera), which introduces shared frozen random matrices across layers to improve efficiency further, and RoseLoRA (wang2024roselora), which employs a row- and column-wise sparse low-rank adaptation mechanism to selectively update the most significant parameters. FourierFT (gaoparameter) replaces the matrix multiplication in LoRA with a Fourier transform, while PiSSA (pissa) and MiLoRA (milora) update the principal and minor singular components of the weight matrix, respectively. However, existing PEFT methods rely on linear transformations to approximate pre-trained weight updates, which struggle to capture the complex relationships inherent in weight updates, leading to a significant performance gap compared to full fine-tuning. Meanwhile, existing research like (teney2024neuralredshift) also demonstrates that nonlinear activation is an integral part of the neural network driving its success.
Figure 1: Framework of proposed PEANuT.
@@ 3 Methodology
In this section, we start with a brief introduction of LoRA. Motivated by a key limitation in LoRA parameter efficiency that roots from LoRA parameterization form, we propose PEANuT, a novel PEFT method to solve the issue. Notably, PEANuT is able to achieves better parameter efficiency provably.
@@ 3.1 Preliminary
LoRA (hu2021lora) assumes that the updates to model weights during the fine-tuning exhibit low-rank properties. Built upon this, LoRA models the incremental update of some weight matrix $\mathbf{W}^{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ in a pre-trained model approximately by the product of two learnable low-rank matrices
@@ 3.2 Inherent Limitation of LoRA Formulation
Specifically, when fully fine-tuning a pre-trained model, the update process of weight $\mathbf{W}$ is typically performed through an iterative gradient descent:
@@
Nonetheless, LoRA matrices $\mathbf{A}$ and $\mathbf{B}$ are parameterized in a free way without any dependency on $\mathbf{W}^{0}$. While gradient $\nabla_{\mathbf{A}}\mathcal{L}$ and $\nabla_{\mathbf{B}}\mathcal{L}$ are implicit functions of $\mathbf{W}^{0}$, making final learned $\mathbf{A}_{t},\mathbf{B}_{t}$ indirectly depends on $\mathbf{W}^{0}$ as well, as will be proved shortly, the lack of explicit dependency still makes LoRA inherently suboptimal for fine-tuning pre-trained models.
3.3 Parameter-Efficient Adaptation with Weight-aware Neural Tweakers
Motivated by the above analysis on LoRA’s limitation, we propose to approximate $\Delta\mathbf{W}$ using a lightweight neural network that explicitly takes pre-trained model weight $\mathbf{W}^{0}$ as input and outputs the weight update directly. By doing so, our approach captures more complex and richer transformation of the weights in a more efficient manner. We refer to our method as parameter-efficient adaptation method with weight-aware neural tweakers (PEANuT).
Following LoRA’s updates paradigm, the proposed PEANuT also provides incremental update of pre-trained models. However, PEANuT modifies the forward pass of the model by introducing a dynamic nonlinear weight transformation. Specifically, the modified model’s forward propagation is formulated as:
@@ 3.3 Parameter-Efficient Adaptation with Weight-aware Neural Tweakers
To ensure the parameter efficiency of our PEANuT, the learnable neural network $f(\mathbf{W}^{0};\bm{\theta})$ should be lightweight, i.e., the number of parameters $\bm{\theta}$ should be much fewer than that of the original pre-trained weight $\mathbf{W}^{0}$. Therefore, we parametrize $f(\mathbf{W}^{0};\bm{\theta})$ as a neural network with bottleneck layers. For example, a simple case is $f(\mathbf{W}^{0};\bm{\theta})=\sigma(\mathbf{W}^{0}\bm{\Theta}_{1})\bm{\Theta}_{2}$, where $\bm{\theta}=(\bm{\Theta}_{1},\bm{\Theta}_{2})\in\mathbb{R}^{d_{2}\times r}\times\mathbb{R}^{r\times d_{2}}$ with $r\ll\min(d_{1},d_{2})$, and $\sigma(\cdot)$ is some non-linear activation function such as ReLU (glorot2011deep). We can also increase the layers or add activation function for the output of $f(\mathbf{W}^{0};\bm{\theta})$ to enhance the model expressiveness.
@@
where the original pre-trained weight $\mathbf{W}^{0}$ is frozen, and only neural network parameters $\theta$ are updated. The overview of PEANuT is shown in Fig. 1.
Remark 3.1.
The benefit of our new formulation lies in two folds. First, our incremental update $f(\mathbf{W}^{0};\bm{\theta})$ is an explicit function of $\mathbf{W}^{0}$, allowing it to capture updates in a more effective way. Second, the neural network-based $f(\mathbf{W}^{0};\bm{\theta})$ allows for dynamic, non-linear weight updates that can capture more complex interactions. These two advantages make PEANuT a more effective and efficient PEFT method than existing LoRA-based approaches.
@@ 3.4 Theoretical Analysis
In this section, we show the theoretical analysis of the sub-optimality of LoRA in terms of parameter efficiency. We prove that PEANuT can achieve equivalent or even superior efficiency under certain conditions. Specifically, suppose PEANuT adopts the following lightweight architecture, as described in Section 3.3:
@@
The following proposition demonstrates that PEANuT can match the expressivity of LoRA using fewer parameters under specific conditions. Here, expressivity is measured by the minimum attainable loss.
@@
In words, Prop 3.2 demonstrates the (approximate) equivalence of LoRA and PEANuT in terms of their expressivity. Specifically, the minimum attainable loss using rank-$r$ LoRA can be achieved by PEANuT with $2r$ hidden units, and conversely, the minimum attainable loss using PEANuT with $r$ hidden units can be achieved rank-$r$ LoRA, provided the invariance assumption holds. This equivalence further implies that the function classes realized by PEANuT with $O(r)$ hidden dimensions and rank-$r$ LoRA are equivalent in expressivity, as the result holds for any loss functions.
Importantly, this highlights a potential improvement in parameter efficiency by PEANuT. Namely, PEANuT with $O(rd_{2})$ parameters maintains the expressivity of LoRA with $r(d_{1}+d_{2})$ parameters. That it to say, PEANuT offers a significant improvement in parameter efficiency when $d_{2}\ll d_{1}$ ( a condition that widely holds for the down projection matrix of transformers fully-connected layers (vaswani2017attention, dosovitskiy2021an) ). In such cases, PEANuT provably achieves better parameter efficiency than LoRA. The added parameter efficiency can also improve sample efficiency by allowing the model to learn representations with the same or fewer data points.
The invariance assumption in Proposition 3.2 pertains to the pre-trained model, and asserts that the later layers of the model depends solely on the task-relevant feature space. Given that we fine-tune a pre-trained model, the later layers are expected to capture this task-relevant feature space, which is described by the left singular space of $\mathbf{W}^{0}$. In practice, since the later layers primarily rely on this pre-trained feature space, the principal directions of the pre-trained weight matrix, represented by its singular vectors, encode most of the useful features for downstream tasks. This makes the loss largely invariant to changes outside this subspace. The proof is available in Appendix B.1.
If we consider a sinusoid activation function $\sigma_{\textnormal{p}}(x)=\sin(2\pi x)$, then stronger result that PEANuT has expressivity (almost) greater than or equal to a LoRA with possibly more parameters can be established without the invariance assumption. We defer the result to the Appendix B.2.
4 Complexity Analysis
In this section, we compare the computational and space complexity of PEANuT and LoRA.
Space Complexity. Because we set the introduced parameters of LoRA and PEANuT to be the same, we only discuss the space complexity of the training in this section. Both LoRA and PEANuT require storing the added parameters and their gradients. PEANuT may incur a slightly higher activation memory during backpropagation due to the extra nonlinearity, but our empirical results (see Sec. 5.4) show that this overhead is minimal and does not affect scalability in practice.
Computational complexity. In terms of per-step computation cost, LoRA computes the residual update as $\mathbf{A}\mathbf{B}x$, which costs $\mathcal{O}(d_{1}r+rd_{2})$ per input vector $x$. PEANuT requires computing $f(\mathbf{W}_{0};\theta)x$. When using the aforementioned example with one-hidden layer and having the same latent dimension as LoRA, the main cost is $\mathcal{O}(d_{1}d_{2}r)$. Although this complexity is higher than LoRA’s, both methods benefit from highly matrix-friendly implementations. In practice, our experiments (see Sec. 5.4) show that the empirical training time per step is comparable. Importantly, during inference, both PEANuT and LoRA allow their update modules to be merged into the original weight matrix $\mathbf{W}_{0}$, ensuring that no additional forward-pass cost is incurred in deployment.